# Machine-Made Media: Monitoring the Mobilization of Machine-Generated Articles on Misinformation and Mainstream News Websites

### 17 Jan 2024 v4

@@ 5 The Rise of Machine-Generated Media
As seen in Table 10, after the release of ChatGPT on November 30, 2022, we observe a noted jump (0.47%) above expectation in the number of synthetic articles from misinformation websites. Every popularity ranking bracket of misinformation websites except for misinformation websites between the ranks of 10K and 100K, saw a statistically significant increase in the absolute percentage of their articles that were synthetic, with misinformation websites in the Rank $>$10M+ popularity bracket seeing the highest jump of 1.52%. This was visually seen in Figure 5. Despite this, we observe that only websites in the popularity bracket 100K $<$ Rank $<$ 1M saw the rate at which the percentage of synthetic articles increases, also increase (i.e., increase in the rate of increase).
We similarly find that each group of mainstream websites with popularity rank $>$1M saw a marked increase in synthetic articles immediately following the release of ChatGPT on November 30, 2022. We further observe a trend increase for several mainstream website popularity brackets. Combined with our misinformation website results, this suggests that smaller, less popular, and otherwise less monitored websites were the ones that saw the biggest increase in synthetic articles following the release of ChatGPT. Indeed, the number of synthetic articles among all groups of websites has been increasing and was at its highest levels on May 1, 2023 (Figure 5). We see this mirrored in the overall increase in the trend of mainstream websites’ use of synthetic articles (increase in the rate of increase) in Table 10. We note that while this analysis is not causal, it illustrates the noticeable increase in the percentage of synthetic articles among misinformation websites immediately following the release of ChatGPT.
@@
All +$0.47\%^{***}$ +$0.006\%^{**}$ +$0.04\%$ +$0.002\%$
Rank $<$ 10K +$0.16\%^{***}$ +$0.003\%$ +$0.04\%$ +$0.003\%^{*}$
10K $<$ Rank $<100$K +$0.18\%$ +$0.01\%$ $0.04\%$ +$0.002\%^{*}$
100K $<$ Rank $<1$M +$0.41\%^{***}$ +$0.006\%^{*}$ +$0.01\%$ +$0.0005\%$
1M $<$ Rank $<10$M +$0.14\%^{***}$ +$0.005\%$ +$0.17\%^{***}$ +$0.002\%$
Rank $>10$M +$1.52\%^{***}$ +$0.005\%$ +$1.10\%^{***}$ +$0.004\%^{***}$
@@ 6 Discussion and Conclusion
In this work, we implement a DeBERTa-based model to classify 15.90 million articles from 3,074 news websites as human-written or synthetic. We find that between January 1, 2022, and May 1, 2023, the percentage of synthetic articles produced by mainstream/reliable news increased by 55.4% while the percentage produced by misinformation/unreliable news websites increased by 457%. Estimating the effect of ChatGPT, we observe a noticeable jump in the percentage of synthetic articles from misinformation websites and unpopular mainstream news around its release. We now discuss several limitations and implications of this work.
Limitations. We note that while we sampled our dataset from a large set of 3,074 news websites and gathered over 15.90M articles, we did not gather articles from every news website and focused on English-language media. As such, our results largely do not apply to non-English media. Similarly, because we used pre-defined lists of misinformation websites, our work largely misses the probable existence of new misinformation websites that appeared since the launch of ChatGPT.
Because we take a conservative approach to our estimation of machine-generated/synthetic texts and due to our removal of articles with characters lengths less than 1000 characters, the absolute numbers presented in this paper are only rough estimates of the percentage of articles on a given website that are machine-generated. As illustrated by Sadasivan et al. (Sadasivan et al. 2023), reliable detection of these short texts is near impossible/largely impractical as large language models become more complex. As shown by Sadasivan et al. (Sadasivan et al. 2023), as LLMs come to more closely match the distribution of written human language, the distinction between human-written and machine-generate texts disappears. As such, we note that while we manage to create a somewhat reliable detector in this work for longer articles for several released and public models, as more advanced and powerful models are developed, effective detection will be more difficult. Similarly, it has been shown that heavily human-edited machine-generated similarly are very difficult to detect as machine-generated (Mitchell et al. 2023) and in this work, we do not seek to detect these instances. As such, due to our conservative approach, our absolute percentage estimates are likely underestimates.
@@
Small Websites and Synthetic Articles. As seen throughout this work, while larger more popular websites have been slower to adopt the use of AI-generate and synthetic content, smaller less popular websites in particular have shown the greatest relative increase in the use of synthetic (728% increase among the least popular misinformation websites and 376% increase among the least popular mainstream websites). We thus find that to fully understand the influence of synthetic media, as similarly argued by News Guard (Sadeghi and Arvanitis 2023), researchers must document and study these less popular websites rather than just concentrating on the top and most frequently visited domains.
The Rise of Synthetic Misinformation. We found that throughout 2022 and 2023, as large language models became more widely accessible, the percentage of machine-generated content on misinformation sites has had a 457% relative increase. While at the beginning of 2022, a lower percentage of misinformation/unreliable news websites’ content was synthetic (0.38% vs. 0.91%), we find that by May 2023, across all popularity brackets examined, misinformation websites had closed this gap (2.14% vs. 1.42%). Unlike popular mainstream websites, misinformation websites and unpopular mainstream websites experienced a noticeable jump in synthetic content after the release of ChatGPT (as determined by our interrupted-time-series analysis). Furthermore, as shown by our topic analysis, misinformation websites have utilized these synthetic articles to address world affairs and health-related news more often than mainstream websites. While not every article posted on an unreliable/misinformation news website is necessarily misinformation, the rapid adoption of synthetic methods by misinformation websites for articles addressing world affairs and health news by these websites could have downstream negative effects. As such given the rapid adoption of the use of synthetic articles by misinformation and unpopular websites, in particular, we argue for future studies of how misinformation websites have utilized these technologies and how the content of these types of articles spread to social media and the broader Internet.
@@ Appendix A Reddit Users and Synthetic Articles
nofilternews 872 dailymail.co.uk 124 investopedia.com 500
autotldr 391 indiatimes.com 15 reuters.com 392
newswall 385 firstpost.com 11 cnbc.com 391
stocks 200 justthenews.com 10 newsweek.com 161
backfieldvacio 142 townhall.com 10 bleacherreport.com 153
@@
Reddit Dataset. To understand Reddit users’ interaction with synthetic content, we first gather all Reddit submissions/posts and their associated metadata (e.g., date posted, the subreddit of the post, number of comments, etc…) that referenced an article published between January 1, 2022, and April 1, 2023, from one of the news websites in our dataset. To collect this Reddit data, we rely upon Pushshift (Baumgartner et al. 2020), which keeps a queryable replica of Reddit data, getting all available submissions posted between January 1, 2022, and March 31, 2023. Altogether, we identify 408,292 Reddit submissions that make use of 486,993 news articles (31,689 from misinformation/unreliable websites; 455,304 from mainstream/reliable news websites) from our list of URLs from 2022 and 2023. Among these articles, we identified 3,470 synthetic articles from mainstream websites and 247 synthetic articles from misinformation websites. We report the subreddits and news websites with the most synthetic articles that appear in our Reddit dataset in Table 11. We note that despite these low numbers, we perform this analysis to get an understanding of the broad trends in interaction with synthetic media on Reddit.
@@
Figure 7: Percentage of Reddit submissions featuring news articles from between January 1, 2022, and March 31, 2023, that went to synthetic/machine-generated articles.
@@
Mainstream Synthetic Articles. As seen in Figure 7, despite the overall relative percentage increase (56.5%) in the use of synthetic new articles by mainstream/reliable news websites, we do not see a correspondingly large percentage increase in the percentage of mainstream/reliable news submissions on Reddit that are synthetic. Between January 1, 2022, and March 31, 2023, we only observed a slight increase from 1.80% to 2.27% (a 25.7% relative increase). However, while synthetic articles in terms of proportions did not increase significantly among posted mainstream articles, in terms of raw numbers, we find that the daily average of mainstream synthetic article submissions went from 4.1 in January 2022 to 7.2 in March 2023 (up 75.6%). Similarly, we observe a raw increase in the average daily number of comments on these submissions from 329.9 comments to 3,134.0 comments, an 850% increase. However, as seen in Figure 8, in terms of the percentage of comments that went to submissions that featured synthetic articles, this corresponded to an absolute percentage decrease of 1.57%.
@@
Misinformation Synthetic Articles. As seen in Figure 7, the percentage of Reddit submissions that feature synthetic news articles from misinformation/unreliable news websites remained relatively stable. We observe only an increase from 2.50% in January 2022, to 2.60% in March 2023, a 3.55% relative increase. However, in terms of raw numbers, we observe that while on average 0.3 synthetic articles were featured in Reddit submissions each day in January 2022, this increased to an average of 0.6 in March 2023, a 100% increase. This corresponds to a similar increase in the number of comments on these submissions. Specifically, we observe a raw increase in the average total number of comments on these submissions each day from 7.2 comments to 13.83 comments, a 92.1% decrease. This corresponded with a 2.92% absolute percentage increase in the percentage of Reddit comments on synthetic articles (Figure 8).

### 20 Mar 2024 v5

@@ 5 The Rise of Machine-Generated Media
As seen in Table 10, after the release of ChatGPT on November 30, 2022, we observe a noted jump (0.50%) above expectation in the number of synthetic articles from misinformation websites. Many of the popularity ranking brackets of misinformation websites saw a statistically significant increase in the absolute percentage of their articles that were synthetic, with misinformation websites in the Rank $>$10M+ popularity bracket seeing the highest jump of 1.68%. This was visually seen in Figure 5. We similarly observe that websites in every popularity bracket except those with Rank $>$10M+ saw the rate at which the percentage of synthetic articles increases, also increase (i.e., increase in the rate of increase).
We further find that the groups of mainstream websites with popularity ranks $>$1M saw a marked increase in synthetic articles immediately following the release of ChatGPT on November 30, 2022. We further observe a trend increase for several mainstream website popularity brackets. Combined with our misinformation website results, this suggests that smaller, less popular, and otherwise less monitored websites were the ones that saw the biggest increase in synthetic articles following the release of ChatGPT. Indeed, the number of synthetic articles among all groups of websites has been increasing and was at its highest levels on May 1, 2023 (Figure 5). We see this mirrored in the overall increase in the trend of mainstream websites’ use of synthetic articles (increase in the rate of increase) in Table 10. We note that while this analysis is not causal, it illustrates the noticeable increase in the percentage of synthetic articles among misinformation websites immediately following the release of ChatGPT.
@@
All +$0.50\%^{***}$ +$0.006\%^{**}$ +$0.04\%$ +$0.001\%$
Rank $<$ 10K +$0.10\%^{***}$ +$0.004\%^{***}$ +$0.03\%$ +$0.003\%^{***}$
10K $<$ Rank $<100$K +$0.10\%$ +$0.01\%$ +$0.03\%$ +$0.001\%$
100K $<$ Rank $<1$M +$0.41\%^{***}$ +$0.007\%^{*}$ +$0.007\%$ +$0.0005\%$
1M $<$ Rank $<10$M +$0.12\%$ +$0.006\%^{*}$ +$0.19\%^{***}$ +$0.002^{*}\%$
Rank $>10$M +$1.68\%^{***}$ +$0.004\%$ +$0.79\%^{***}$ +$0.004\%^{***}$
@@ 6 Discussion and Conclusion
In this work, we implement a DeBERTa-based model to classify 15.46 million articles from 3,074 news websites as human-written or synthetic. We find that between January 1, 2022, and May 1, 2023, the percentage of synthetic articles produced by mainstream/reliable news increased by 57.3% while the percentage produced by misinformation/unreliable news websites increased by 474%. Estimating the effect of ChatGPT, we observe a noticeable jump in the percentage of synthetic articles from misinformation websites and unpopular mainstream news around its release. We now discuss several limitations and implications of this work.
Limitations. We note that while we sampled our dataset from a large set of 3,074 news websites and gathered over 15.46M articles, we did not gather articles from every news website and focused on English-language media. As such, our results largely do not apply to non-English media. Similarly, because we used pre-defined lists of misinformation websites, our work largely misses the probable existence of new misinformation websites that appeared since the launch of ChatGPT.
Because we take a conservative approach to our estimation of machine-generated/synthetic texts and due to our removal of articles with characters lengths less than 1000 characters, the absolute numbers presented in this paper are only rough estimates of the percentage of articles on a given website that are machine-generated. As illustrated by Sadasivan et al. (2023), reliable detection of these short texts is near impossible/largely impractical as large language models become more complex. As shown by Sadasivan et al. (2023), as LLMs come to more closely match the distribution of written human language, the distinction between human-written and machine-generate texts disappears. As such, we note that while we manage to create a somewhat reliable detector in this work for longer articles for several released and public models, as more advanced and powerful models are developed, effective detection will be more difficult. Similarly, it has been shown that heavily human-edited machine-generated similarly are very difficult to detect as machine-generated (Mitchell et al. 2023) and in this work, we do not seek to detect these instances. As such, due to our conservative approach, our absolute percentage estimates are likely underestimates.
@@
Small Websites and Synthetic Articles. As seen throughout this work, while larger more popular websites have been slower to adopt the use of AI-generated and synthetic content, smaller less popular websites in particular have shown the greatest relative increase in the use of synthetic (736% increase among the least popular misinformation websites and 423% increase among the least popular mainstream websites). We thus find that to fully understand the influence of synthetic media, as similarly argued by News Guard (Sadeghi and Arvanitis 2023), researchers must document and study these less popular websites rather than just concentrating on the top and most frequently visited domains.
The Rise of Synthetic Misinformation. We found that throughout 2022 and 2023, as large language models became more widely accessible, the percentage of machine-generated content on misinformation sites has had a 474% relative increase. While at the beginning of 2022, a lower percentage of misinformation/unreliable news websites’ content was synthetic (0.39% vs. 0.88%), we find that by May 2023, across all popularity brackets examined, misinformation websites had closed this gap (2.22% vs. 1.39%). Unlike popular mainstream websites, misinformation websites and unpopular mainstream websites experienced a noticeable jump in synthetic content after the release of ChatGPT (as determined by our interrupted-time-series analysis). Furthermore, as shown by our topic analysis, misinformation websites have utilized these synthetic articles to address world affairs and health-related news more often than mainstream websites. While not every article posted on an unreliable/misinformation news website is necessarily misinformation, the rapid adoption of synthetic methods by misinformation websites for articles addressing world affairs and health news by these websites could have downstream negative effects. As such given the rapid adoption of the use of synthetic articles by misinformation and unpopular websites, in particular, we argue for future studies of how misinformation websites have utilized these technologies and how the content of these types of articles spread to social media and the broader Internet.
@@ Appendix A Reddit Users and Synthetic Articles
nofilternews 984 dailymail.co.uk 96 investopedia.com 500
autotldr 948 globaltimes.cn 34 reuters.com 392
newswall 442 justthenews.com 10 cnbc.com 391
stocks 201 washingtonexaminer.com 9 newsweek.com 161
backfieldvacio 142 express.co.uk 7 bleacherreport.com 153
@@
Reddit Dataset. To understand Reddit users’ interaction with synthetic content, we first gather all Reddit submissions/posts and their associated metadata (e.g., date posted, the subreddit of the post, number of comments, etc…) that referenced an article published between January 1, 2022, and April 1, 2023, from one of the news websites in our dataset. To collect this Reddit data, we rely upon Pushshift (Baumgartner et al. 2020), which keeps a queryable replica of Reddit data, getting all available submissions posted between January 1, 2022, and March 31, 2023. Altogether, we identify a set of Reddit submissions that make use of 254,302 news articles (19,299 from misinformation/unreliable websites; 235,003 from mainstream/reliable news websites) from our list of URLs from 2022 and 2023. Among these articles, we identified 3,428 synthetic articles from mainstream websites and 209 synthetic articles from misinformation websites. We report the subreddits and news websites with the most synthetic articles that appear in our Reddit dataset in Table 11. We note that despite these low numbers, we perform this analysis to get an understanding of the broad trends in interaction with synthetic media on Reddit.
@@
Figure 7: Percentage of Reddit submissions featuring news articles published between January 1, 2022, and March 31, 2023, that went to synthetic/machine-generated articles.
@@
Mainstream Synthetic Articles. As seen in Figure 7, despite the overall relative percentage increase (57.3%) in the use of synthetic new articles by mainstream/reliable news websites, we do not see a correspondingly large percentage increase in the percentage of mainstream/reliable news submissions on Reddit that are synthetic. Between January 1, 2022, and March 31, 2023, we only observed a slight increase from 1.26% to 1.43% (a 12.9% relative increase). However, while synthetic articles in terms of proportions did not increase significantly among posted mainstream articles, in terms of raw numbers, we find that the daily average of mainstream synthetic article submissions went from 4.1 in January 2022 to 7.3 in March 2023 (78.0% increase). Similarly, we observe a raw increase in the average daily number of comments on these submissions from 329.9 comments to 3,134.0 comments, an 850% increase. However, as seen in Figure 8, in terms of the percentage of comments that went to submissions that featured synthetic articles, this corresponded to an absolute percentage decrease of 1.57% (this amount has varied dramatically however).
@@
Misinformation Synthetic Articles. As seen in Figure 7, the percentage of Reddit submissions that feature synthetic news articles from misinformation/unreliable news websites remained relatively stable. We observe only an increase from 2.50% in February 2022 (there was a temporary uptick in January 2022), to 1.67% in March 2023, a 6.39% relative increase. However, in terms of raw numbers, we observe that while on average 0.4 synthetic articles were featured in Reddit submissions each day in February 2022, this increased to an average of 0.5 in March 2023, a 20% increase. This corresponds to a similar increase in the number of comments on these submissions. Specifically, we observe a raw increase in the average total number of comments on these submissions each day from 9.60 comments to 13.67 comments, a 42.6% increase (Figure 8).